OpenStudy (anonymous):

Use integration in parts to integrate [e^x(sinx)]dx. I made u=sin(x), du=cos(x)dx, v=e^x and dv=e^xdx. I solved for the whole thing and everything cancelled out, leaving me with C. Is this correct?

jimthompson5910 (jim_thompson5910):

so you're saying you got \[\Large \int e^x\sin(x)dx = C\] ??

jimthompson5910 (jim_thompson5910):

if so, then it is incorrect

OpenStudy (anonymous):

Bah! I had a feeling it was wrong.

OpenStudy (anonymous):

That is the right equation though, yes.

OpenStudy (anonymous):

is this one of those where you go around in a circle and get the same thing back again ? then divide by 2?

jimthompson5910 (jim_thompson5910):

hopefully you have this so far? \[\Large \int udv = uv - \int v du\] \[\Large \int e^x\sin(x)dx = e^x\sin(x) - \int e^x \cos(x)dx\]

OpenStudy (anonymous):

Yeah, that's what I got halfway through. Am I supposed to stop there or did I integrate wrong?

jimthompson5910 (jim_thompson5910):

you have to use one more iteration of integration by parts

OpenStudy (anonymous):

Ooooh. OH. I should've realized. WELP, let's see if I can make this work!

jimthompson5910 (jim_thompson5910):

u = sin(x), du = cos(x)dx dv = e^x, v = e^x \[\large \int udv = uv - \int v du\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - \int e^x \cos(x)dx\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - \color{red}{\int e^x \cos(x)dx}\] w = cos(x), dw = -sin(x)dx dz = e^x, z = e^x \[\large \int e^x\sin(x)dx = e^x\sin(x) - \color{red}{[wz - \int z dw]}\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - \color{red}{[e^x\cos(x) - \int e^x(-\sin(x)dx)]}\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - \color{black}{[e^x\cos(x) + \int e^x\sin(x)dx]}\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - e^x\cos(x) - \int e^x\sin(x)dx\]

OpenStudy (anonymous):

Yep, just worked out that part! And when I solve out for the last part of the equation, will I end up with e^xsin(x) - 2(e^xcos(x))?

jimthompson5910 (jim_thompson5910):

very close

jimthompson5910 (jim_thompson5910):

but no

OpenStudy (anonymous):

Ah, wait, that would just cancel out again.

jimthompson5910 (jim_thompson5910):

Let \[M = \int e^x\sin(x)dx\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - e^x\cos(x) - \int e^x\sin(x)dx\] \[\large M = e^x\sin(x) - e^x\cos(x) - M\] solve for M and re-substitute

OpenStudy (anonymous):

Is it e^xsin(x) - e^xcos(x) + C?

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

A right Ouroboros of pain and trig, this problem. Okay. What next?

jimthompson5910 (jim_thompson5910):

Let \[M = \int e^x\sin(x)dx\] \[\large \int e^x\sin(x)dx = e^x\sin(x) - e^x\cos(x) - \int e^x\sin(x)dx\] \[\large M = e^x\sin(x) - e^x\cos(x) - M\] \[\large M+M = e^x\sin(x) - e^x\cos(x) - M+M\] \[\large 2M = e^x\sin(x) - e^x\cos(x)\] \[\large M = \frac{e^x\sin(x) - e^x\cos(x)}{2}+C\] \[\large \int e^x\sin(x)dx = \frac{e^x\sin(x) - e^x\cos(x)}{2}+C\]

OpenStudy (anonymous):

Bless you, kind sir.

jimthompson5910 (jim_thompson5910):

no problem